OpenStudy (anonymous):

Show that \[\int_0^\infty \frac{dx}{1+x^n}=\frac{\pi}{n}\csc\frac{\pi}{n}\] for \(n\in\mathbb{N}\backslash\{1\}\). Edit: another relationship that appears to be true: \[\int_0^\infty \frac{dx}{1+x^n}=\frac{1}{n}\Gamma\left(\frac{1}{n}\right)\Gamma\left(1-\frac{1}{n}\right)=\frac{1}{n}\mathrm{B}\left(\frac{1}{n},1-\frac{1}{n}\right)\]

OpenStudy (anonymous):

The identities \[\Gamma(x)\Gamma(-x)=-\frac{\pi}{x\sin(\pi x)}\quad\text{and}\quad\Gamma(x)\Gamma(1-x)=\frac{\pi}{\sin (\pi x)}\] seem to shed some light on the relation above. (Identities 41 and 42 shown here: http://mathworld.wolfram.com/GammaFunction.html )

OpenStudy (anonymous):

Applying the residue theorem and using a half-circle contour over the upper-half plane seems an okay strategy, but calculating all those residues... Not sure if induction would help, but then again, I have yet to try.

OpenStudy (freckles):

\[\int\limits_{0}^{\infty}\frac{dx}{1+x^2}=\arctan(x)|_0^{\infty}=\frac{\pi}{2}=\frac{\pi}{2} \csc(\frac{\pi}{2})\] So it works for n=2 now assume it works for some integer k that is we have \[\int\limits_{0}^{\infty}\frac{dx}{1+x^k}=\frac{\pi}{k} \csc(\frac{\pi}{k})\] now let's show this is true for k+1 that is we want to show \[\int\limits_{0}^{\infty}\frac{dx}{1+x^{k+1}}=\frac{\pi}{k+1} \csc(\frac{\pi}{k+1})\] hmm.... I wonder if is possible to prove it by induction actually stuck right now here

OpenStudy (anonymous):

great job

OpenStudy (anonymous):

Here's something we might be able to use... For \(|x|<1\), \[\frac{1}{1+x^n}=\frac{1}{1-(-x^n)}=\sum_{k=0}^\infty(-1)^kx^{nk}\] For \(|x|>1\), \[\frac{1}{1+x^n}=\frac{1}{x^n}\frac{1}{1-(-x^{-n})}=\frac{1}{x^n}\sum_{k=0}^\infty (-1)^kx^{-nk}=\sum_{k=0}^\infty (-1)^k x^{-n(k+1)}\] which we know converges because the ratio test gives us this result: \[\lim_{k\to\infty}\left|\frac{(-1)^{k+1} x^{-n(k+2)}}{(-1)^k x^{-n(k+1)}}\right|=|x|^{-n}<1~~\iff~~1<|x|^n~~\implies~~1<|x|\] So, \[\begin{align*} \int_0^\infty \frac{dx}{1+x^n} &=\left\{\int_0^1+\int_1^\infty\right\}\frac{dx}{1+x^n}\\\\ &=\sum_{k=0}^\infty \left\{\int_0^1 (-1)^kx^{nk}\,dx+\int_1^\infty (-1)^kx^{-n(k+1)}\,dx\right\}\\\\ &=\sum_{k=0}^\infty \left\{(-1)^k\left[\frac{x^{nk+1}}{nk+1}\right]_{0}^{1}+(-1)^k\left[\frac{x^{-n(k+1)+1}}{-n(k+1)+1}\right]_1^\infty\right\}\\\\ &=\sum_{k=0}^\infty \left\{\frac{(-1)^k}{nk+1}-\frac{(-1)^{k}}{1-n(k+1)}\right\}\\\\ &=\left(1-\frac{1}{1-n}\right)+\left(-\frac{1}{n+1}+\frac{1}{1-2n}\right)\\ &\quad\quad+\left(\frac{1}{2n+1}-\frac{1}{1-3n}\right)+\left(-\frac{1}{3n+1}+\frac{1}{1-4n}\right)+\cdots\\\\ &=1-\frac{2}{1-n^2}+\frac{2}{1-4n^2}-\frac{2}{1-9n^2}+\cdots\\\\ &=\sum_{k=0}^\infty \frac{(-2)^k}{1-(kn)^2} \end{align*}\] Hmm, this kinda resembles the Basel problem.

OpenStudy (anonymous):

Actually, there might be a flaw in this work somewhere along the line, since \[\lim_{k\to\infty}\left|\frac{(-2)^{k+1}}{1-(k+1)^2n^2}\times\frac{1-k^2n^2}{(-2)^k}\right|=2\lim_{k\to\infty}\left|\frac{\frac{1}{k^2}-n^2}{1-\left(\frac{k+1}{k}\right)^2n^2}\right|=2>1\] and so the series appears to diverge by the ratio test...

OpenStudy (anonymous):

Apparently, the contour method seems to work out nicely when you use a contour that surrounds only one of the complex roots, as shown here: http://math.stackexchange.com/questions/247866/show-that-int-0-infty-frac11xn-dx-frac-pi-n-sin-pi-n-wh But I think I'm more interested in a real method of approaching this.

ganeshie8 (ganeshie8):

sub \(u = \dfrac{1}{1+x^n} \implies x = \left(\dfrac{1}{u}-1\right)^{1/n}\) the integral becomes : \[\begin{align}\int_0^\infty \frac{dx}{1+x^n} &= \frac{1}{n} \int_0^1 u^{-\frac{1}{n}}(1-u)^{\frac{1}{n}-1} \,du\\~\\ &=\frac{1}{n} B\left(1-\frac{1}{n}, ~\frac{1}{n}\right)\\~\\ &= \frac{1}{n}\dfrac{\Gamma(1-\frac{1}{n}) \Gamma(\frac{1}{n})}{\Gamma(1)}\\~\\ &=\frac{1}{n} \frac{\pi}{\sin(\frac{\pi}{n})} \end{align}\] for a more general case see http://openstudy.com/users/ganeshie8#/updates/54a7145ce4b054f0c3b7e558

ganeshie8 (ganeshie8):

induction idea seems new xD

OpenStudy (anonymous):

Oh yeah I remember seeing that question, totally slipped my mind